OpenStudy (luigi0210):

Simple derivative function:

OpenStudy (luigi0210):

\[y=(sect+tant)^5\]

terenzreignz (terenzreignz):

Chain rule again. Once again, think of this: \[\Large y = (\color{blue}{\sec t + \tan t })^5\] as just one entity, differentiate as per the power rule... and then multiply the whole thing by the derivative of that 'entity'

OpenStudy (psymon):

Multiplication of the derivatives of layers (inner functions).

OpenStudy (ybarrap):

$$ \bf \frac{d}{dx}(\sec(x) + \tan(x))^5\\\\ \text{Let u= }(\sec(x) + \tan(x))\\\\ \text{Then }\frac{du^5}{du}\frac{du}{dx}=5u^4\frac{du}{dx}\\\\ \frac{du}{dx}=\frac{d}{dx}(\sec(x) + \tan(x))=\sec^2(x))+\sec(x)\tan(x)\\\\ \text{So, }\frac{d}{dx}(\sec(x) + \tan(x))^5\\\\ =5(\sec(x) + \tan(x))^4(\sec^2(x))+\sec(x)\tan(x))\\\\ $$